OpenStudy (anonymous):

Solve, using the substitution method. y–2x=8 16+4x=2y A. There are an infinite number of solutions. B. The only solution is (1, 10). C. The only solution is (24, 0). D. There is no solution.

OpenStudy (jdoe0001):

if you solve the 1st equation for "y", what do you get?

OpenStudy (anonymous):

Y-2x=8?

OpenStudy (jdoe0001):

well, we know that much, so.. and solving for "y" gives us?

OpenStudy (anonymous):

y-2x= 8 y-2x+2x=8+2x y=2x+8

OpenStudy (anonymous):

Y=2x+8 .... 16+4x=2(2x+8) .. 16+4x=4x+16 so infinity ... so a? @jdoe0001

OpenStudy (jdoe0001):

was.... doing something but yes, correct